OpenStudy (anonymous):

How do I work out 5cos(squared)x + 4sinxcosx + 2sin(squared)x = 2 ???

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

the equation simplifies to 3cos^2(x) +4sin(x)cos(x)=0 , take cos(x0 out and hence u get cos(x)(3cos(x) +4sin(x))=0 , => cos(x)=0 or tan(x)=-4/3 , use these to getto the solution. note tan(53 degrees)=4/3.

OpenStudy (anonymous):

sorry tan(x)=-3/4 is wat u get...tan(37 degrees) is =3/4 :)